OpenStudy (anonymous):

Raphael is twice as old as his sister is. The sum of their ages is less than 30. What is the greatest age Raphael’s sister can be?

OpenStudy (shinalcantara):

let 'x' be the age of his sister 'y' be Raphael's age y = 2x ----equation 1 x + y < 30 ----inequality 1 substitute equation 1 to the inequality x + 2x < 30 3x < 30 x < 10 if it her age be less than 10 then the oldest she could be is x-1 x - 1 = 10 -1 = 9 yrs old